OpenStudy (anonymous):

The height h in feet of a baseball on Earth after t seconds can be modeled by the function h(t) = -16(t – 1.5)2 + 36, where -16 is a constant in ft/s2 due to Earth's gravity. The gravity on Mars is only 0.38 times that on Earth. If the same baseball were thrown on Mars, it would reach a maximum height 59 feet higher and 2.5 seconds later than on Earth. Write a height function for the baseball thrown on Mars.

OpenStudy (zale101):

Label earth's gravity, height, and time based on the given equation \(h(t) = -16(t – 1.5)^2 + 36\)

OpenStudy (zale101):

What is -16 ? they've already told you is a constant in ft/s2 due to Earth's gravity. That means, -16 ft/s^2 is the gravity

OpenStudy (zale101):

First, let's look closely on what the graph supposed to represent. h(t) = -16(t – 1.5)2 + 36 f(t)=a(x-h)^2+k (h,k) is the vertex We know from an upside down parabola, the vertex is at the maximum height. h (x axis) represent the time in which the objects change Height on k ( y-axis) 

OpenStudy (zale101):

Clear @glosan1010 ?

OpenStudy (zale101):

f(x)=a(x-h)^2+k Gravity represents a time represents h height represents k The projectile motion on earth is h(t) = -16(t – 1.5)^2 + 36 Gravity is...? Time is...? and Height is...?

OpenStudy (zale101):

If you know your gravity, time, and height, you can easily use the earth's projectile motion equation and convert it to the projectile motion on mars.